# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ III-G Optimizing The TBS MAs’ Position
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow{\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n]\}$ (57)
$-\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n]\}+c_{4,3,m,n}\leq 0,$
where $c_{4,3,m,n}$ represents a constant term and the above newly introduced coefficients defined in (58).
$\mathbf{F}^{10}_{k}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),\mathbf{F}^{11}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),\mathbf{F}^{12}\triangleq({\sum}_{k=1}^{K_{d}}\mathbf{F}^{10}_{k}+\mathbf{F}^{11})/\Gamma,$ (58)
$\mathbf{F}^{13}_{k,j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),\mathbf{F}^{14}_{j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),\mathbf{F}^{15}_{j}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{F}^{13}_{k,j}+\mathbf{F}^{14}_{j},$
$\bar{\mathbf{F}}^{15}_{j,m,n}\triangleq\mathbf{F}^{15}_{j}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],f^{12}_{j,m,n}\triangleq{\sum}_{z\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{15}_{j,m,m}[z,n])^{\ast}\mathbf{g}_{t,m,j}[z],$
$\bar{\mathbf{F}}^{12}_{m,n}\triangleq\mathbf{F}^{12}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],{f}^{15}_{m,n}\triangleq{\sum}_{i\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{12}_{m,m}[i,n])^{\ast}\mathbf{g}_{t,m,0}[i],$
$\mathbf{f}^{11}_{j,m}\triangleq{\sum}_{i\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{15}_{j,i,m})^{H}\mathbf{g}_{t,i,j},\mathbf{f}^{14}_{m}\triangleq{\sum}_{n\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{12}_{m,n})^{H}\mathbf{g}_{t,n,0},f^{13}_{j,m,n}\triangleq f^{12}_{j,m,n}+\mathbf{f}^{11}_{j,m}[n],{f}^{16}_{m,n}\triangleq{f}^{15}_{m,n}+\mathbf{f}^{14}_{m}[n].$
Based on the above equivalent transformation, the subproblem w.r.t $\mathbf{t}^{0}_{m,n}$ is formulated as
[…]
$+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})\}\big{)}+\hat{c}_{4,2,m,k,n}$
$+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}$ (59a)
s.t. ${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}$ (59b)
$-\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}+c_{4,20,m,n}\leq 0,$
$\|\mathbf{t}^{0}_{m,n}-\mathbf{t}^{0}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (59c)
$\mathbf{t}^{0}_{m,n}\in\mathcal{C},$ (59d)
@@
where $\hat{c}_{4,2,m,k,n}$ is a constant. Obviously, the current form of problem (P9) is difficult to obtain the optimal $\mathbf{t}^{0}_{m,n}$ due to the non-convex objective function and constraints. In the following, we adopt the MM methodology [20] to resolve the above difficulty.
Firstly, for any positive semidefinite matrix $\mathbf{F}$, we have the following inequality
$\mathbf{h}^{H}\mathbf{F}\mathbf{h}$ (60)
@@
with $\lambda_{max}(\mathbf{F})$ being the maximal eigenvalue of $\mathbf{F}$, $c_{F}$ is a constant, $\mathbf{h}_{0}\triangleq\mathbf{h}^{0}_{m,k,n,0}(\mathbf{t}^{0}_{m,n,0})$, $\mathbf{t}^{0}_{m,n,0}=[x^{0}_{m,n,0},y^{0}_{m,n,0}]^{T}$ is the latest value of $\mathbf{t}^{0}_{m,n}$. Next, by leveraging the arguments presented in (60), we can construct a tight upper-bound for $(\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n}))^{H}\mathbf{F}_{m,k,n}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})$, which can be formulated in (61), where $\|\mathbf{h}^{0}_{m,k,n}\|^{2}_{2}=\|\mathbf{h}^{0}_{m,k,n,0}\|^{2}_{2}=L^{t,0}_{m,k}$, $c_{4,4,m,k,n}\triangleq 2\lambda_{max}(\mathbf{F}_{5,m,k,n,n})L^{t,0}_{m,k}-((\mathbf{h}^{0}_{m,k,n,0})^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n,0})^{\ast}$, $\mathbf{f}^{17}_{m,k,n}\!\triangleq\!2((\mathbf{F}_{5,m,k,n,n})^{H}\mathbf{h}^{0}_{m,k,n,0}\!-\!\lambda_{max}(\mathbf{F}_{5,m,k,n,n})\mathbf{h}^{0}_{m,k,n,0})$ and $\mathbf{f}^{18}_{m,k,n}\!\triangleq\!\mathbf{f}^{17}_{m,k,n}\!+\!\mathbf{f}^{7}_{m,k,n}$.
$(\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n}))^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})\}$ (61)
$\leq\lambda_{max}(\mathbf{F}_{5,m,k,n,n})\|\mathbf{h}^{0}_{m,k,n}\|^{2}_{2}-2\textrm{Re}\{\lambda_{max}(\!\mathbf{F}_{5,m,k,n,n}\!)(\mathbf{h}^{0}_{m,k,n,0})^{H}\mathbf{h}^{0}_{m,k,n}\}\!+\!\!\lambda_{max}(\!\mathbf{F}_{5,m,k,n,n}\!)\|\mathbf{h}^{0}_{m,k,n,0}\|^{2}_{2}$
+2Re {($+\!\!2\textrm{Re }\!\!\{(\mathbf{h}^{0}_{m,k,n,0})^{H}\!\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n}\}-\!\!2\textrm{Re }\!\!\{(\mathbf{h}^{0}_{m,k,n,0})^{H}\!\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n,0}\}\!\!+\!\!(\mathbf{h}^{0}_{m,k,n,0})^{H}\mathbf{F}_{5,m,k,n,n}\mathbf{h}^{0}_{m,k,n,0}$
$=\textrm{Re}\{(\mathbf{f}^{17}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}+c_{4,21,m,k,n}+\textrm{Re}\{(\mathbf{f}^{7}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}(\mathbf{t}^{0}_{m,n})\}=\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}+c_{4,4,m,k,n}$
Next, to tackle the challenging terms $\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}$, $\textrm{Re}\{({f}^{10}_{m,j,n})^{\ast}\mathbf{g}_{t,m,j}[n](\mathbf{t}^{0}_{m,n})\}$, $\textrm{Re}\{(f^{13}_{i,m,n})^{\ast}\mathbf{g}_{t,m,i}[n](\mathbf{t}^{0}_{m,n})\}$ and $\textrm{Re}\{(f^{16}_{m,n})^{\ast}\mathbf{g}_{t,m,0}[n](\mathbf{t}^{0}_{m,n})\}$, we introduce the following results that are proved in [3].
@@
Assume that $\mathbf{f}\in\mathbb{C}^{L\times 1}$, $\mathbf{h}(\mathbf{t})\in\mathbb{C}^{L\times 1}$ and $\mathbf{t}=[x,y]^{T}$. The convex lower and upper bounds of the function $h_{t}(\mathbf{t})\triangleq\textrm{Re}\{\mathbf{f}^{H}\mathbf{h}(\mathbf{t})\}=\sum_{l=1}^{L}|\mathbf{f}[l]||\mathbf{h}(\mathbf{t})[l]|\cos(\angle\mathbf{h}(\mathbf{t})[l]-\angle\mathbf{f}[l])$ can be written as
$h_{t}(\mathbf{t})\!\geq\!h_{t}(\mathbf{t}_{0})\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!-\!\frac{\delta}{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (62)
$h_{t}(\mathbf{t})\!\leq\!h_{t}(\mathbf{t}_{0})\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!+\!\frac{\delta}{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (63)
@@
respectively, where $\mathbf{t}_{0}$ is obtained from the last iteration, the symbols |||\ || | and $\angle$ denote the operations of obtaining the amplitude and phase, respectively, $\delta$ is a positive real number satisfying $\delta=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L}|\mathbf{f}[l]|$, $\nabla h_{t}(\mathbf{t})=\big{[}\frac{\partial h_{t}(\mathbf{t})}{\partial x},\frac{\partial h_{t}(\mathbf{t})}{\partial y}\big{]}^{T}\in\mathbb{C}^{2\times 1}$ denotes the gradient vector of $h_{t}(\mathbf{t})$ over $\mathbf{t}$.
@@
Firstly, by using the inequality (63) in Lemma 1, we can obtain convex upper-bound of $\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}$ as shown in (64),
$\textrm{Re}\{(\mathbf{f}^{18}_{m,k,n})^{H}\mathbf{h}^{0}_{m,k,n}\}={\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{18}_{m,k,n}[l]|\cos\big{(}\frac{2\pi}{\lambda}(\mathbf{t}^{0}_{m,n})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{18}_{m,k,n}[l]\big{)}\triangleq h^{t0}_{,m,k,n}(\mathbf{t}^{0}_{m,n})$ (64)
$\leq h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})+\nabla^{T}h^{t0}_{m,k,n}(\mathbf{t}^{0}_{m,n,0})(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,n})+\frac{\delta_{1,m,k,n}}{2}(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,n})^{T}(\mathbf{t}^{0}_{m,n,0}-\mathbf{t}^{0}_{m,n})$
$=\frac{\delta_{1,m,k,n}}{2}\|\mathbf{t}^{0}_{m,n}\|^{2}_{2}+(\mathbf{f}^{19}_{m,k,n})^{T}\mathbf{t}^{0}_{m,n}+c_{4,5,m,k,n}$

### v3

@@ III-G Optimizing The TBS MAs’ Position
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow$ (59)
${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{4}_{i})^{\ast}{g}_{t,i}\}-\textrm{Re}\{(f^{5})^{\ast}\mathbf{g}_{t,0}\}+c_{4,3}\leq 0,$
where $c_{4,3}$ is constant and the newly introduced coefficients are defined as follows
$\mathbf{F}^{10}_{k}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),$ (60)
$\mathbf{F}^{11}\triangleq\sigma_{t,0}^{2}(\hat{\mathbf{g}}_{r,0}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),$
$\mathbf{F}^{13}_{k,j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}),$
$\mathbf{F}^{14}_{j}\triangleq\sigma_{t,j}^{2}(\hat{\mathbf{g}}_{r,j}^{H}\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j})^{T}\otimes(\hat{\mathbf{W}}_{r}\hat{\mathbf{W}}_{r}^{H}),$
$\mathbf{F}^{12}\triangleq({\sum}_{k=1}^{K_{d}}\mathbf{F}^{10}_{k}+\mathbf{F}^{11})/\Gamma,\mathbf{F}^{15}_{j}\triangleq{\sum}_{k=1}^{K_{d}}\mathbf{F}^{13}_{k,j}+\mathbf{F}^{14}_{j},$
$\bar{\mathbf{F}}^{15}_{j,m,n}\triangleq\mathbf{F}^{15}_{j}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],$
$f^{22}_{j,m,n}\triangleq{\sum}_{z\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{15}_{j,m,m}[z,n])^{\ast}\mathbf{g}_{t,m,j}[z],$
$\bar{\mathbf{F}}^{12}_{m,n}\triangleq\mathbf{F}^{12}[(m-1)N_{t}+1:mN_{t},(n-1)N_{t}+1:nN_{t}],$
${f}^{24}_{m,n}\triangleq{\sum}_{i\neq n}^{N_{t}}2(\bar{\mathbf{F}}^{12}_{m,m}[i,n])^{\ast}\mathbf{g}_{t,m,0}[i],$
$\mathbf{f}^{21}_{j,m}\triangleq{\sum}_{i\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{15}_{j,i,m})^{H}\mathbf{g}_{t,i,j},\mathbf{f}^{23}_{m}\triangleq{\sum}_{n\neq m}^{M_{t}}2(\bar{\mathbf{F}}^{12}_{m,n})^{H}\mathbf{g}_{t,n,0},$
$f^{4}_{j}\triangleq f^{22}_{j,m,n}+\mathbf{f}^{21}_{j,m}[n],{f}^{5}\triangleq{f}^{24}_{m,n}+\mathbf{f}^{23}_{m}[n].$
Based on the above equivalent transformation, the subproblem w.r.t $\mathbf{t}^{0}$ is formulated as
$\textrm{(P9)}:$ $\mathop{\textrm{min}}\limits_{\mathbf{t}^{0}}{\sum}_{k=1}^{K_{d}}\big((\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})$
$+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}-\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}\big)$
$+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}+{c}_{4,4}$ (61a)
s.t. ${\sum}_{i=1}^{K_{t}}\textrm{Re}\{(f^{4}_{i})^{\ast}{g}_{t,i}(\mathbf{t}^{0})\}$ (61b)
$-\textrm{Re}\{(f^{5})^{\ast}{g}_{t,0}(\mathbf{t}^{0})\}+c_{4,3}\leq 0,$
$\|\mathbf{t}^{0}-\mathbf{t}^{0}_{m,i}\|_{2}\geq D_{t},\forall n,i\in\mathcal{N}_{t},n\neq i,$ (61c)
$\mathbf{t}^{0}\in\mathcal{C},$ (61d)
@@
where ${c}_{4,4}$ is a constant. Obviously, the current form of problem (P9) is difficult to obtain the optimal $\mathbf{t}^{0}$ due to the non-convex objective function and constraints. In the following, we adopt the MM methodology [35] to resolve the above difficulty. Firstly, for any positive semidefinite matrix $\mathbf{F}$, we have the following inequality
$\mathbf{h}^{H}(\mathbf{t}^{0})\mathbf{F}\mathbf{h}(\mathbf{t}^{0})$ (62)
@@
with $\lambda_{max}(\mathbf{F})$ being the maximal eigenvalue of the matrix $\mathbf{F}$, $c_{F}$ is a constant, $\mathbf{h}_{0}\triangleq\mathbf{h}(\mathbf{t}^{0}_{0})$, $\mathbf{t}^{0}_{0}=[x^{0}_{0},y^{0}_{0}]^{T}$ is the latest value of $\mathbf{t}^{0}$.
Next, by leveraging the arguments presented in (62), we can construct a tight upper-bound for $(\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})$, which can be formulated in (63), where $\mathbf{h}^{0}_{k,0}\triangleq\mathbf{h}^{0}_{k}(\mathbf{t}^{0}_{0})$, $\|\mathbf{h}^{0}_{k}\|^{2}_{2}=\|\mathbf{h}^{0}_{k,0}\|^{2}_{2}=L^{t,0}_{m,k}$, $c_{4,5,k}\triangleq 2\lambda_{max}(\mathbf{{F}}^{1}_{k})L^{t,0}_{m,k}-((\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k,0})^{\ast}$, $\mathbf{f}^{6}_{k}\!\triangleq\!2\big((\mathbf{{F}}^{1}_{k})^{H}\mathbf{h}^{0}_{k,0}\!-\!\lambda_{max}(\mathbf{{F}}^{1}_{k})\mathbf{h}^{0}_{k,0}\big)$ and $\mathbf{f}^{7}_{k}\!\triangleq\!\mathbf{f}^{6}_{k}\!+\!\mathbf{f}^{1}_{k}$.
$(\mathbf{h}^{0}_{k}(\mathbf{t}^{0}))^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}$ (63)
$\leq\lambda_{max}(\mathbf{{F}}^{1}_{k})\|\mathbf{h}^{0}_{k}\|^{2}_{2}-2\textrm{Re}\{\lambda_{max}(\mathbf{{F}}^{1}_{k})(\mathbf{h}^{0}_{k,0})^{H}\mathbf{h}^{0}_{k}\}+\lambda_{max}(\mathbf{{F}}^{1}_{k})\|\mathbf{h}^{0}_{k,0}\|^{2}_{2}$
$+2\textrm{Re }\{(\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k}\}-2\textrm{Re }\{(\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k,0}\}+(\mathbf{h}^{0}_{k,0})^{H}\mathbf{{F}}^{1}_{k}\mathbf{h}^{0}_{k,0}+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}$
$=\textrm{Re}\{(\mathbf{f}^{6}_{k})^{H}\mathbf{h}^{0}_{k}\}+c_{4,5,k}+\textrm{Re}\{(\mathbf{f}^{1}_{k})^{H}\mathbf{h}^{0}_{k}(\mathbf{t}^{0})\}=\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}+c_{4,5,k}$
Furthermore, to convexity the challenging terms $\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}$, $\textrm{Re}\{(\mathbf{f}^{2}_{k})^{H}\mathbf{h}^{0}_{k}\}$ and $\textrm{Re}\{({f}^{3}_{j})^{\ast}{g}_{t,j}(\mathbf{t}^{0})\}$, we introduce the following results that are proved in [4].
@@
Assume that $\mathbf{f}\in\mathbb{C}^{L\times 1}$, $\mathbf{h}(\mathbf{t})=[e^{j\!\frac{2\pi}{\lambda}\!(\cos(\!\theta_{1}\!)\sin(\!\phi_{1}\!)x\!+\!\sin(\theta_{1})y)},\cdots,e^{j\!\frac{2\pi}{\lambda}\!(\cos(\!\theta_{L}\!)\sin(\!\phi_{L}\!)x\!+\!\sin(\theta_{L})y)}]\in\mathbb{C}^{L\times 1}$ and $\mathbf{t}=[x,y]^{T}$. The convex lower and upper bounds of the function $h_{t}(\mathbf{t})\triangleq\textrm{Re}\{\mathbf{f}^{H}\mathbf{h}(\mathbf{t})\}=\sum_{l=1}^{L}|\mathbf{f}[l]|\cos(\angle\mathbf{h}(\mathbf{t})[l]-\angle\mathbf{f}[l])$ ( the variable $\mathbf{t}$ only exists in the phase part of $h_{t}(\mathbf{t})$ ) can be written as
$h_{t}(\mathbf{t})\!\geq\!h_{t}(\mathbf{t}_{0})\!\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!-\!{\delta}/{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (64)
$h_{t}(\mathbf{t})\!\leq\!h_{t}(\mathbf{t}_{0})\!\!+\!\nabla h_{t}^{T}(\mathbf{t}_{0})(\mathbf{t}\!-\!\mathbf{t}_{0})\!+\!{\delta}/{2}(\mathbf{t}\!-\!\mathbf{t}_{0})^{T}(\mathbf{t}\!-\!\mathbf{t}_{0}),$ (65)
@@
respectively, where $\mathbf{t}_{0}$ is obtained from the last iteration, the symbols |||\ | and $\angle$ denote the operations of obtaining the amplitude and phase, respectively, $\delta$ is a positive real number satisfying $\delta=\frac{8\pi^{2}}{\lambda^{2}}{\sum}_{l=1}^{L}|\mathbf{f}[l]|$, $\nabla h_{t}(\mathbf{t})=\big[\frac{\partial h_{t}(\mathbf{t})}{\partial x},\frac{\partial h_{t}(\mathbf{t})}{\partial y}\big]^{T}\in\mathbb{C}^{2\times 1}$ denotes the gradient vector of $h_{t}(\mathbf{t})$ over $\mathbf{t}$.
@@
Firstly, by using the inequality (65) in Lemma 1, we can obtain a convex upper-bound of $\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}$, showing as
$\textrm{Re}\{(\mathbf{f}^{7}_{k})^{H}\mathbf{h}^{0}_{k}\}\triangleq h^{1}_{k}(\mathbf{t}^{0})$ (66)
$={\sum}_{l=1}^{L^{t,0}_{m,k}}|\mathbf{f}^{7}_{k}[l]|\cos\big(({2\pi}/{\lambda})(\mathbf{t}^{0})^{T}\mathbf{a}^{t,0}_{m,k,l}-\angle\mathbf{f}^{7}_{k}[l]\big)$
$\leq h^{1}_{k}(\mathbf{t}^{0}_{0})+\nabla^{T}h^{1}_{k}(\mathbf{t}^{0}_{0})(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})+0.5{\delta_{1,k}}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})^{T}(\mathbf{t}^{0}-\mathbf{t}^{0}_{0})$
$=0.5{\delta_{1,k}}\|\mathbf{t}^{0}\|^{2}_{2}+(\mathbf{f}^{8}_{k})^{T}\mathbf{t}^{0}+c_{4,6,k},$